OpenStudy (rissyroo118):

18x^3-30x^2 ______________ 6x 2x^3 2x^4 3x^2-5x 3x^4-5x medal goes to whoever gives me the best answer

OpenStudy (jdoe0001):

is that a question or a statement?

OpenStudy (rissyroo118):

question

TheSmartOne (thesmartone):

This is your question: \(\sf\Large\frac{18x^3-30x^2}{6x}\) Use this to simplify it. \(\sf\Large\frac{a-b}{c}=\frac{a}{c}-\frac{b}{c}\)

OpenStudy (michele_laino):

hint: \[\frac{{18{x^3} - 30{x^2}}}{{6x}} = \frac{{6x\left( {3{x^2} - 5x} \right)}}{{6x}} = ...?\]